# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ IV Proofs for Problem Formulation
We prove Propositions 1, 3, 2 and 1 in this section. Throughout this proof, we employ a concise notation by using overlapping labels (e.g., $c,\tilde{c},c_{1},c_{2},c^{\prime},\hat{c},\dots$) to denote various constants. To avoid ambiguity, we explicitly reserve the notation $c_{\alpha},c_{\beta},c_{\gamma}$ for specially defined constants in Lemma 7 and the subsequent proof.
@@ IV-A Proof of Proposition 1
It is easy to verify that $\pi\triangleq\big{(}\sum_{i=1}^{n}\alpha_{i}\big{)}^{-1}[\alpha_{1},\dots,\alpha_{n}\big{]}^{\mathrm{T}}$ is the stationary distribution of the canonical Markov matrix $S$ corresponding to $P$. Moreover, since $S$ is reversible, therefore for any $(i,j)\in{\mathcal{E}}$ we have
@@
For the converse, since $p_{ij}>0$ for all $(i,j)\in{\mathcal{E}}$, $S$ exhibits irreducibility since the graph ${\mathcal{G}}$ is strongly-connected, and aperiodicity as $S_{ii}>0$. As a result, $S$ has a unique stationary distribution $\pi$. By reversibility of $S$, we have for all $i\neq j$ and $(i,j)\in{\mathcal{E}}$,
@@
where last step follows from the fact that $p_{ij}+p_{ji}=1$. Thus, $P$ corresponds to a BTL model with weight vector $\pi$.
@@ IV-B Proof of Proposition 2
and the stationary distribution satisfies $\Pi P+P\Pi={\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})$. Also, the stationary distribution is independent of graph structure (as long as the adjacency matrix of graph is symmetric). Conversely, if $\Pi P+P\Pi={\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})$ is true, this implies that
@@
This completes the proof.
@@ IV-C Proof of Proposition 3
$\|\Pi P+P\Pi-$ ${\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|^{2}_{\pi^{-1},\mathrm{F}}=\|\Pi P-P^{\mathrm{T}}\Pi\|^{2}_{\pi^{-1},\mathrm{F}}+\|(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\Pi\|^{2}_{\pi^{-1},\mathrm{F}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +2\text{Tr}\left(\Pi^{-1/2}(\Pi P-P^{\mathrm{T}}\Pi)^{\mathrm{T}}(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\Pi^{1/2}\right)$
$=\|\Pi P-P^{\mathrm{T}}\Pi\|^{2}_{\pi^{-1},\mathrm{F}}+\|P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}})\|^{2}_{\pi,\mathrm{F}}+2\text{Tr}\left((\Pi P-P\Pi)^{\mathrm{T}}(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\right).$
Now it remains to show that $\text{Tr}\left((\Pi P-P^{\mathrm{T}}\Pi)^{\mathrm{T}}(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\right)=0$. This follows from the fact that $\text{Tr}(A^{\mathrm{T}}B)=0$ when $A$ is anti-symmetric, i.e., $A^{\mathrm{T}}=-A$, and $B$ is symmetric, i.e., $B=B^{\mathrm{T}}$:
$\text{Tr}(A^{\mathrm{T}}B)$ $=-\text{Tr}(AB)=-\text{Tr}(AB^{\mathrm{T}})=-\text{Tr}(A^{\mathrm{T}}B)\implies\text{Tr}(A^{\mathrm{T}}B)=0.$
Clearly $\Pi P-P^{\mathrm{T}}\Pi$ is anti-symmetric and the symmetry of $P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}})$ follows since the set ${\mathcal{E}}$ is symmetric by assumption. This completes the proof.
@@ IV-D Proof of Theorem 1
We begin our proof by establishing the upper bound. Assume that for a BTL model $B\in\mathsf{BTL}_{h}$, its skill score parameters are given by $\pi_{B}\in{\mathbb{R}}^{n}_{+}$. We will interchangeably use the notation $\pi_{B}\in\mathsf{BTL}_{h}$ to mean $B\in\mathsf{BTL}_{h}$. Moreover, without loss of generality, we assume that $\sum_{i=1}^{n}\pi_{B,i}=1$, where $\pi_{B,i}$ is the $i$th component of vector $\pi_{B}$ and moreover, let $\Pi_{B}\triangleq\operatorname{{diag}}(\pi_{B})$ and $\pi_{\min}\triangleq\min_{i\in[n]}\pi_{i}$. Thus, we have
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $=\min_{\pi_{B}\in\mathsf{BTL}_{h}}\sqrt{\sum_{(i,j)\in{\mathcal{E}}}\frac{((\pi_{B,i}+\pi_{B,j})p_{ij}-\pi_{B,j})^{2}}{(\pi_{B,i}+\pi_{B,j})^{2}}}$
@@
where $\pi_{B,\min}\triangleq\min_{i\in[n]}\pi_{B,i}$, $\zeta_{1}$ follows from the fact that $\pi_{B,\min}\geq\|\pi_{B}\|_{\infty}/h$ as $B\in\mathsf{BTL}_{h}$, and $\zeta_{2}$ follows by substituting $\pi_{B}=\pi$ which are the skill score parameters obtained from the stationary distribution of the canonical Markov matrix $S$ corresponding to $P$ and by assumption $\pi$ belongs to the set $\mathsf{BTL}_{h}$.
Now, we will focus on proving the lower bound. Specifically, we will show that there exist constants $c,c^{\prime}$ such that
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $\geq c\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\|_{\mathrm{F}},$ (30)
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $\geq c^{\prime}\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{-1/2}P^{\mathrm{T}}\Pi^{1/2}\|_{\mathrm{F}}.$ (31)
If both Eqs. 30 and 31 are true, then we can use the following argument to complete the proof. Let $c_{0}=\min\{c,c^{\prime}\}$, then we have:
$2\min_{B\in\mathsf{BTL}_{h}}\|P-B\|_{\mathrm{F}}$ $\geq c_{0}(\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{-1/2}P^{\mathrm{T}}\Pi^{1/2}\|_{\mathrm{F}}+\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\|_{\mathrm{F}})$ (32)
$\stackrel{{\scriptstyle\zeta_{1}}}{{\geq}}c_{0}\bigg{(}\frac{\|\Pi P-P^{\mathrm{T}}\Pi\|_{\pi^{-1},\mathrm{F}}}{\sqrt{\|\pi\|_{\infty}}}+\frac{\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\|_{\pi,\mathrm{F}}}{\sqrt{\|\pi\|_{\infty}}}\bigg{)}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\geq}}c_{0}\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\pi^{-1},\mathrm{F}}}{\sqrt{\|\pi\|_{\infty}}}$
$\geq c_{0}\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}}{{{\|\pi\|_{\infty}}}},$
where $\zeta_{1}$ follows since $\|AB\|_{\mathrm{F}}\geq\sigma_{n}(A)\|B\|_{\mathrm{F}}$ and $\zeta_{2}$ follows from Proposition 3 and sub-additivity of square-root.
@@
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $\geq\min_{\begin{subarray}{c}B\in[0,1]^{n\times n}:\\ B+B^{\mathrm{T}}={\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}\end{subarray}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}=\frac{1}{2}\left\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\right\|_{\mathrm{F}},$
@@
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $\geq\min_{B\in\mathsf{BTL}_{h}}\sqrt{\frac{\pi_{B,\min}}{\|\pi_{B}\|_{\infty}}}\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}-\Pi_{B}^{1/2}{\mathcal{P}_{\mathcal{E}}}(B)\Pi_{B}^{-1/2}\|_{\mathrm{F}}$
=$\stackrel{{\scriptstyle\zeta_{1}}}{{=}}\min_{B\in\mathsf{BTL}_{h}}\sqrt{\frac{\pi_{B,\min}}{\|\pi_{B}\|_{\infty}}}\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}-{\mathcal{P}_{\mathcal{E}}}(\Pi_{B}^{1/2}B\Pi_{B}^{-1/2})\|_{\mathrm{F}}$
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}-\Pi_{B}^{-1/2}P^{\mathrm{T}}\Pi_{B}^{1/2}\|_{\mathrm{F}}\geq c\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{1/2}P^{\mathrm{T}}\Pi^{-1/2}\|_{\mathrm{F}}.$
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|_{\mathrm{F}}\geq\tilde{c}\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}.$ (33)
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}$ $-\Pi_{B}^{-1/2}P^{\mathrm{T}}\Pi_{B}^{1/2}\|_{\mathrm{F}}\geq\min_{\pi_{B}\in\mathsf{BTL}_{h}}\frac{\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|_{\mathrm{F}}}{\|\pi_{B}\|_{\infty}}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\geq}}\min_{\pi_{B}\in\mathsf{BTL}_{h}}\frac{n}{h}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|_{\mathrm{F}}\stackrel{{\scriptstyle\zeta_{2}}}{{\geq}}\frac{n\tilde{c}}{h}\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}$
$\geq\frac{\tilde{c}n\pi_{\min}}{h}\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{-1/2}P^{\mathrm{T}}\Pi^{1/2}\|_{\mathrm{F}}$
$\stackrel{{\scriptstyle\zeta_{3}}}{{\geq}}\frac{\tilde{c}}{h^{2}}\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{-1/2}P^{\mathrm{T}}\Pi^{1/2}\|_{\mathrm{F}},$
where $\zeta_{1}$ follows since $\pi_{B}\in\mathsf{BTL}_{h}$, hence $\pi_{B,i}\geq\|\pi_{B}\|_{\infty}/h$ which implies $\|\pi_{B}\|_{\infty}\leq\frac{h}{n}$, $\zeta_{2}$ follows from our assumption in Eq. 33 (which we will prove next) and $\zeta_{3}$ follows since $\pi_{\min}\geq\|\pi\|_{\infty}/h$ and $n\|\pi\|_{\infty}\geq 1$.
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|^{2}_{\mathrm{F}}\geq\min_{\pi_{B}\in{\mathbb{R}}^{n}:\pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}=1}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|^{2}_{\mathrm{F}}.$
@@
$\min_{\pi_{B}\in{\mathbb{R}}^{n}:\pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}=1}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|^{2}_{\mathrm{F}},$
and let ${\Pi_{B}^{*}}\triangleq\operatorname{{diag}}({\pi_{B}^{*}})$. To prove Eq. 33, we will show that the following ratio is upper bounded by a constant.
$\frac{\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}}{\|{\Pi_{B}^{*}}P-P^{\mathrm{T}}{\Pi_{B}^{*}}\|_{\mathrm{F}}}=\sqrt{\frac{\pi^{\mathrm{T}}R\pi}{\pi_{B}^{*\mathrm{T}}R\pi_{B}^{*}}}.$ (34)

### v2

@@ IV Proofs for Problem Formulation
We prove Propositions 1, 3, 2 and 1 in this section. Throughout this and future sections, we employ a concise notation by using overlapping labels (e.g., $c,\tilde{c},c_{1},c_{2},c^{\prime},\hat{c}$$\dots$) to denote various constants. To avoid ambiguity, we explicitly reserve the notation $c_{\alpha},c_{\beta},c_{\gamma}$ for specially defined constants in Lemma 7 and the subsequent proof (these special constants can depend on parameters like $n$). We also note that some of the bounds used in the subsequent analysis (in this and future sections) are valid only with high probability. For brevity, we do not restate this at every occurrence.
@@ IV-A Proof of Proposition 1
It is easy to verify that $\pi\triangleq\big(\sum_{i=1}^{n}\alpha_{i}\big)^{-1}[\alpha_{1},\dots,\alpha_{n}\big]^{\mathrm{T}}$ is the stationary distribution of the canonical Markov matrix $S$ corresponding to $P$. Moreover, since $S$ is reversible, for any $(i,j)\in{\mathcal{E}}$, we have
@@
For the converse, since $p_{ij}>0$ for all $(i,j)\in{\mathcal{E}}$, $S$ exhibits irreducibility since the graph $\mathcal{G}$} is strongly-connected, and aperiodicity as $S_{ii}>0$. As a result, $S$ has a unique stationary distribution $\pi$. By reversibility of $S$, we have for all $(i,j)\in{\mathcal{E}}$ with $i\neq j$,
@@
where last step follows from the fact that $p_{ij}+p_{ji}=1$. Thus, $P$ corresponds to a BTL model with parameters $\pi$. ∎
@@ IV-B Proof of Proposition 2
and the stationary distribution satisfies $\Pi P+P\Pi={\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})$. Also, the stationary distribution is independent of graph structure (as long as the adjacency matrix of the graph is symmetric). Conversely, if $\Pi P+P\Pi={\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})$ is true, this implies that
@@
This completes the proof. ∎
@@ IV-C Proof of Proposition 3
$\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|^{2}_{\pi^{-1},\mathrm{F}}=\|\Pi P-P^{\mathrm{T}}\Pi\|^{2}_{\pi^{-1},\mathrm{F}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\|(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\Pi\|^{2}_{\pi^{-1},\mathrm{F}}+$
$2\text{Tr}\left(\Pi^{-1/2}(\Pi P-P^{\mathrm{T}}\Pi)^{\mathrm{T}}(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\Pi^{1/2}\right)$
$=\|\Pi P-P^{\mathrm{T}}\Pi\|^{2}_{\pi^{-1},\mathrm{F}}+\|P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}})\|^{2}_{\pi,\mathrm{F}}+$
$\ \ \ \ \ \ \ \ \ \ \ \ \ 2\text{Tr}\left((\Pi P-P^{\mathrm{T}}\Pi)^{\mathrm{T}}(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\right).$
Now it remains to show that $\text{Tr}\big((\Pi P-P^{\mathrm{T}}\Pi)^{\mathrm{T}}(P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}}))\big)=0$. This follows from the fact that $\text{Tr}(A^{\mathrm{T}}B)=0$ when $A$ is anti-symmetric, i.e., $A^{\mathrm{T}}=-A$, and $B$ is symmetric, i.e., $B=B^{\mathrm{T}}$:
$\text{Tr}(A^{\mathrm{T}}B)$ $=-\text{Tr}(AB)=-\text{Tr}(AB^{\mathrm{T}})=-\text{Tr}(A^{\mathrm{T}}B)$
$\implies\text{Tr}(A^{\mathrm{T}}B)=0.$
Clearly, $\Pi P-P^{\mathrm{T}}\Pi$ is anti-symmetric, and the symmetry of $P^{\mathrm{T}}+P-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\mathbf{1}_{n}^{\mathrm{T}})$ follows since the set $\mathcal{E}$} is symmetric by assumption. This completes the proof. ∎
@@ IV-D Proof of Theorem 1
We begin our proof by establishing the upper bound. Assume that for a BTL model $B\in\mathsf{BTL}_{h}$, its skill score parameters are given by $\pi_{B}\in{\mathbb{R}}^{n}_{+}$. We will interchangeably use the notation $\pi_{B}\in\mathsf{BTL}_{h}$ to mean $B\in\mathsf{BTL}_{h}$. Moreover, without loss of generality, we assume that $\sum_{i=1}^{n}\pi_{B,i}=1$, where $\pi_{B,i}$ is the $i$th component of vector $\pi_{B}$ and moreover, let $\Pi_{B}\triangleq\operatorname{{diag}}(\pi_{B})$. Thus, we have
$\min_{B\in\mathsf{BTL}_{h}}$ $\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$
$=\min_{\pi_{B}\in\mathsf{BTL}_{h}}\sqrt{\sum_{(i,j)\in{\mathcal{E}}}\frac{((\pi_{B,i}+\pi_{B,j})p_{ij}-\pi_{B,j})^{2}}{(\pi_{B,i}+\pi_{B,j})^{2}}}$
@@
where $\pi_{B,\min}\triangleq\min_{i\in[n]}\pi_{B,i}$, $\zeta_{1}$ follows from the fact that $\pi_{B,\min}\geq\|\pi_{B}\|_{\infty}/h$ as $B\in\mathsf{BTL}_{h}$, and $\zeta_{2}$ follows by substituting $\pi_{B}=\pi$, which are the skill score parameters obtained from the stationary distribution of the canonical Markov matrix $S$ corresponding to $P$, and by Assumption 2, $\pi$ belongs to the set $\mathsf{BTL}_{h}$.
Now, we will focus on proving the lower bound. Specifically, we will show that there exists a constant $c^{\prime}$ such that
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $\geq\frac{1}{2}\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\|_{\mathrm{F}},$ (28)
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $\geq c^{\prime}\frac{\xi^{2}}{h^{3}}\|\Pi^{\frac{1}{2}}P\Pi^{-\frac{1}{2}}\!-\!\Pi^{-\frac{1}{2}}P^{\mathrm{T}}\Pi^{\frac{1}{2}}\|_{\mathrm{F}}.$ (29)
If both Eqs. 28 and 29 are true, then we can use the following argument to complete the proof. Letting $c_{1}=\min\{0.5,c^{\prime}\frac{\xi^{2}}{h^{3}}\}$, we have
$2\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$
$\geq c_{1}(\|\Pi^{\frac{1}{2}}P\Pi^{-\frac{1}{2}}\!-\!\Pi^{-\frac{1}{2}}P^{\mathrm{T}}\Pi^{\frac{1}{2}}\|_{\mathrm{F}}\!+\!\|P+\!P^{\mathrm{T}}\!-\!\!{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\|_{\mathrm{F}})$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\geq}}c_{1}\bigg(\frac{\|\Pi P-P^{\mathrm{T}}\Pi\|_{\pi^{-1},\mathrm{F}}}{\sqrt{\|\pi\|_{\infty}}}+\frac{\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\|_{\pi,\mathrm{F}}}{\sqrt{\|\pi\|_{\infty}}}\bigg)$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\geq}}c_{1}\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\pi^{-1},\mathrm{F}}}{\sqrt{\|\pi\|_{\infty}}}$
$\geq c_{1}\frac{\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}}{{{\|\pi\|_{\infty}}}},$ (30)
where $\zeta_{1}$ follows since $\|AB\|_{\mathrm{F}}\geq\sigma_{n}(A)\|B\|_{\mathrm{F}}$ and $\zeta_{2}$ follows from Proposition 3 and the sub-additivity of the square root function.
@@
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$ $\geq\min_{\begin{subarray}{c}B\in[0,1]^{n\times n}:\\ B+B^{\mathrm{T}}={\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}\end{subarray}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$
$=\frac{1}{2}\left\|P+P^{\mathrm{T}}-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}})\right\|_{\mathrm{F}},$
@@
$\min_{B\in\mathsf{BTL}_{h}}\|P-{\mathcal{P}_{\mathcal{E}}}(B)\|_{\mathrm{F}}$
$\geq\min_{B\in\mathsf{BTL}_{h}}\sqrt{\frac{\pi_{B,\min}}{\|\pi_{B}\|_{\infty}}}\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}-\Pi_{B}^{1/2}{\mathcal{P}_{\mathcal{E}}}(B)\Pi_{B}^{-1/2}\|_{\mathrm{F}}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{=}}\min_{B\in\mathsf{BTL}_{h}}\sqrt{\frac{\pi_{B,\min}}{\|\pi_{B}\|_{\infty}}}\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}-{\mathcal{P}_{\mathcal{E}}}(\Pi_{B}^{1/2}B\Pi_{B}^{-1/2})\|_{\mathrm{F}}$
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}$ $\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}-\Pi_{B}^{-1/2}P^{\mathrm{T}}\Pi_{B}^{1/2}\|_{\mathrm{F}}$
$\ \ \ \ \ \ \ \ \ \geq\frac{c\xi^{2}}{h^{5/2}}\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{1/2}P^{\mathrm{T}}\Pi^{-1/2}\|_{\mathrm{F}}.$
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|_{\mathrm{F}}\geq\frac{\tilde{c}\xi^{2}}{h^{1/2}}\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}.$ (31)
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}$ $\|\Pi_{B}^{1/2}P\Pi_{B}^{-1/2}-\Pi_{B}^{-1/2}P^{\mathrm{T}}\Pi_{B}^{1/2}\|_{\mathrm{F}}$
$\geq\min_{\pi_{B}\in\mathsf{BTL}_{h}}\frac{\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|_{\mathrm{F}}}{\|\pi_{B}\|_{\infty}}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\geq}}\min_{\pi_{B}\in\mathsf{BTL}_{h}}\frac{n}{h}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|_{\mathrm{F}}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\geq}}\frac{\tilde{c}n\xi^{2}}{h^{3/2}}\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\geq}}\frac{\tilde{c}n\xi^{2}\pi_{\min}}{h^{3/2}}\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{-1/2}P^{\mathrm{T}}\Pi^{1/2}\|_{\mathrm{F}}$
$\stackrel{{\scriptstyle\zeta_{3}}}{{\geq}}\frac{\tilde{c}\xi^{2}}{h^{5/2}}\|\Pi^{1/2}P\Pi^{-1/2}-\Pi^{-1/2}P^{\mathrm{T}}\Pi^{1/2}\|_{\mathrm{F}},$
where $\zeta_{1}$ follows since $\pi_{B}\in\mathsf{BTL}_{h}$, hence $\pi_{B,i}\geq\|\pi_{B}\|_{\infty}/h$ which implies that $\|\pi_{B}\|_{\infty}\leq\frac{h}{n}$, $\zeta_{2}$ follows from our assumption in Eq. 31 (which we will prove next) and we define $\pi_{\min}\triangleq\min_{i\in[n]}\pi_{i}$, and $\zeta_{3}$ follows since $\pi_{\min}\geq\|\pi\|_{\infty}/h$ and $n\|\pi\|_{\infty}\geq 1$.
@@
$\min_{\pi_{B}\in\mathsf{BTL}_{h}}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|^{2}_{\mathrm{F}}\geq\min_{\begin{subarray}{c}\pi_{B}\in{\mathbb{R}}_{+}^{n}:\\ \pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}=1\end{subarray}}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|^{2}_{\mathrm{F}}.$
@@
$\min_{\pi_{B}\in{\mathbb{R}}_{+}^{n}:\pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}=1}\|\Pi_{B}P-P^{\mathrm{T}}\Pi_{B}\|^{2}_{\mathrm{F}},$
and let ${\Pi_{B}^{*}}\triangleq\operatorname{{diag}}({\pi_{B}^{*}})$. To prove Eq. 31, we will show that the following ratio is upper bounded by $O(\sqrt{h}/\xi^{2})$:
$\frac{\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}}{\|{\Pi_{B}^{*}}P-P^{\mathrm{T}}{\Pi_{B}^{*}}\|_{\mathrm{F}}}=\sqrt{\frac{\pi^{\mathrm{T}}R\pi}{\pi_{B}^{*\mathrm{T}}R\pi_{B}^{*}}},$ (32)